# WixUp: A General Data Augmentation Framework for Wireless Perception in Tracking of Humans

### v1

@@ 3 Background: FMCW raw data processing and formats
So, we briefly introduce the standard FMCW-based data processing pipeline in the majority of wireless perception systems, including the standard mmWave radar [32], and many acoustic sensing and WiFi sensing systems adopted similar radar modulation [35, 29, 62, 61]. Fig. 2 shows the pipeline including de-chirping and extraction of range, angle, velocity, and then 3D Cartesian coordinates. In detail, firstly, the transmitter (TX) sends out chirps; by reflecting from the subject, it is captured by the receiver (RX). Then the signals are processed through a mixer and filter to produce a mixed Intermediate Frequency (IF) signal (This could also achieved by algorithms outside the hardware when a mixer is not available). This signal undergoes a Fast Fourier Transform (FFT) process, where the amplitudes can induce range measurement. Subsequently, another Doppler-FFT is applied to the phases along the slow time for velocity measurement. The multi-channel phases can further estimate the angles of arrival. Then a Constant False Alarm Rate (CFAR) algorithm filtered the outputs by the noise level. Finally, the ranges and angles can be transformed to Cartesian coordinates along with velocity and signal intensity as 5D times series for downstream applications. This pipeline enables detecting position and movement in human tracking or other applications.
@@
The challenge is the prevalent lack of raw data availability within the majority of public datasets, while the raw data, rich in detail, is critically important for performing meaningful and effective data augmentation. These datasets typically provide only Cartesian coordinates, i.e. the culmination of the data processing sequence. The issue with Cartesian coordinates is their high-level abstraction, owing to the filtering by CFAR, which leads to a considerable loss of fine-grained information that is imperative for in-depth data augmentation. Moreover, the sparsity issue in the processed point clouds also makes it worse, while the range profile is a relatively dense representation. Unfortunately, it is intractable to reverse the CFAR algorithmically like inverse FFT.
@@
Thereby, we aim to propose a framework that can transform any data format of wireless signals into range profiles. Because the range profile is lossless, unlike Cartesian coordinates, and it is more interpretable compared with the raw data, allowing in-depth DA. The following subsections will detail how we achieve this goal, focusing on developing a way to inverse Cartesian coordinates to a simulated range profile.
4 Method
Figure 3: A demonstration of the data mixing pipeline of WixUp.
As the high-level illustration shown in Fig. 3, WixUp uses a common process of mixing-based DA. So, in this section, we first provide a formal definition of the mix operation in the context of wireless.
@@ 4 Method
Next, we detail WixUp’s algorithm, tailored for mixing wireless signals at the range-profile level, and how it is embedded into the common mixing-based process. In overview, as in Fig 3, we first describe our solution to transform coordinates into range profiles using Gaussian mixtures. Next, we define a fault-tolerant mixing method leveraging the intersection of the Gaussians, with only O(n) computation complexity. Additionally, we use bootstrapping to further densify the mixed output. Finally, if the downstream model input is not a range profile but coordinates, we induce 3D coordinates from the intersected range profiles along with the angles based on probability distribution.
Furthermore, since we aim to test WixUp’s capability of self-learning for reducing labeling efforts, this section also briefly introduces the process of self-training to instruct our experiment design.
4.1 Define mixing-based data augmentation
@@
A typical mixing-based DA has the inputs involving two sets of data: let $m$ and $n$ denote the sensing data and their ground truth. The input for a mixing operation is a pair of data $(m_{0},n_{0})$ and $(m_{1},n_{1})$. The output is synthesized data, $(m_{0}^{\prime},n_{0}^{\prime})$, in the same format as the two input tuples. Therefore, by iteratively mixing pairs of $(m_{i},n_{i})$ and $(m_{i+1},n_{i+1})$ where $i=0,1,...l$ ($l$ is the size of the original data), we could yield $l-1$ of synthesized data by one pass of iteration. Moreover, the distance $d$ between each pair can vary from 1 to 2 or more, i.e. $(m_{i},n_{i})$ and $(m_{i+d},n_{i+d})$. Thereby, by iterating with $d=1,2,...s$, we scale up the data by $s$ times.
@@
What DA focuses on is how we mix $(m_{0},n_{0})$ and $(m_{1},n_{1})$, especially $m_{0}$ and $m_{1}$. To elaborate, $n$ is usually mixed by a convex combination, i.e. weighted average. In our paper, we take the average regardless of task-specific ground truth format. For example, in pose estimation, $n$ could be a $19\times 3$ matrix, representing the 3D skeleton of 19 body keypoints’ coordinates of $(x,y,z)$. Taking the average of $n_{0}$ and $n_{1}$ equals to the 3D skeleton located in the middle between the two input skeletons. In user identification or action recognition tasks, the $n$ could be the class ID and is usually one-hot encoded. Therefore, taking the average of two one-hot vectors means the mixed data is half-chance class A and half-chance class B, or the same class when the input classes are equivalent. As for the data $m$, in the DA for 2D images, they could be a matrix of pixels in RGB channels, where a popular mixing-based DA undergoes taking convex combination in the same way of $n$. However, in wireless perception, $m$ could be raw signals, range profiles, 3D point clouds, or 5D time series of $(x,y,z,D,I)$ where $D$ is the Doppler velocity and $I$ is the signal intensity. As discussed above, directly adding them in some way does not meaningfully align with the feature of wireless signals. Instead, WixUp aims to transform them all into the range profile level for the convenience of mixing with rich raw information. Next, we will introduce how we mix $m_{0}$ and $m_{1}$ by a pipeline of signal processing and probability-based algorithms.
4.2 WixUp data mixing pipeline
@@
4.2.1 Inverse Cartesian coordinates to simulated range profile using Gaussian mixture
@@
It is possible for most states in the FMCW data processing pipeline to freely transform to range profile using the standard signal processing algorithms. Because most of the steps can be reverse-engineered, such as inverse FFT, so that we can transform any data format into a range profile. However, the CFAR presents a unique challenge as it inherently discards a significant quantity of data, making reversal complex. Therefore, we put forth a novel inverse data processing pipeline that operates between the range data and Cartesian coordinates. This enables the application of WixUp to any public datasets, irrespective of the data format they are provided in, allowing us to bypass the limitation imposed by the absence of raw data.
@@
The core idea is to simulate the range profile as a Gaussian mixture model of Euclidean distances, as shown in the left of Fig. 3. In detail, first, we translate Cartesian coordinates to spherical coordinates as a representation better aligned with the transmission features of wireless signals. Next, by dividing the Euclidean range by range resolution, we get indices of bins. Utilizing these indices as the means in a Gaussian mixture model, the probability density function (PDF) is derived as the range profile, i.e. the statistical distribution of the ranges. The parameters defined for the system include a window size of 512, which is the size for de-chirping data, and a range resolution calculated as 3.75cm. Then the maximum allowable detection range is 19.2m. Assuming the human joint size is around 10-15cm, we use the standard Gaussian, implying its main lobe falls within the joint size.
4.2.2 Mixing range profiles using intersection
@@
As illustrated in the center of Fig. 3, each pair of peaks may intersect midway along the range axis; the height of the intersected side lobe is roughly inversely proportional to the distance between the peaks. This approach involves taking middle points from all possible bisections of two sets of real points, as it is unclear which body joint each real point corresponds to; thus, we cannot restrict middle points to be extracted from identical joints. By considering all possible bisections, we ensure coverage of eligible synthetic points and introduce some "temperature" into the process, borrowed from sampling in VAE. This makes the mixing fault-tolerant.
Instead of employing the closed-form intersection of Gaussian algorithms, we utilize an O(n) one-pass function to identify intersections. The validity of intersections can be determined using the formula $(b_{i}-a_{i})(b_{i-1}-a_{i-1})<0$, where $a$ and $b$ are the two PDF arrays and $(a_{i},a_{i-1})$ denotes a neighbor pair. This approach ensures that $ab$ pairs like (0,0)(0,0) and (0,1)(0,0) are excluded. This exclusion is crucial as such pairs do not represent valid intersection points. By implementing this method, we effectively filter out non-intersecting segments by one pass. This streamlines the intersection detection process, enhancing both efficiency and accuracy and does not rely on the complexity of the input Gaussian mixture.
4.2.3 Boostrapping high-probability synthesis
Besides, we further bootstrap each intersection point weighting on its height. Since the height of the intersection is proportional to its validity as an eligible point, we use it as a weight to resample points around it. This follows the Monte Carlo simulation, where random samples are generated to simulate the properties of a system or to approximate integrals and sums. Besides, to add more randomness, we also keep the original peaks in the simulated range profile. They are assigned the lowest weight of one and all the weights of the above intersections are elevated by one as well.
@@ 4.2.3 Boostrapping high-probability synthesis
4.2.4 Enrich range profile with angles based on probability distribution double-paned
@@
4.3 Use WixUp for unsupervised domain adaptation via self-training
We not only aim to evaluate our data augmentation in standard supervised learning for wireless perception tasks, but also demonstrate its capability for reducing labeling efforts via self-training. This is achieved through our method of data mixing to generate new data. Note that non-mixing methods like random scaling are not a good fit, since they augment one data at a time, unlike mixing two data from different domains to close the domain gap. Mixing-based augmentation enables the creation of more realistic semi-real labels by mixing real data and predicted data, facilitating efficient domain adaptation through self-training.
Source domain & Target domain: The process of training and evaluating unsupervised domain adaptation using self-training is to adapt models trained on a source domain to inference well on a target domain. In this approach, the target domain dataset might have scarce labeled data or even no labeled data. In step 1 of this approach, the model is first trained with labeled data from the source domain. In step 2, partial target domain data is split out as training data and the rest is for testing. If we make predictions on testing target data using the trained model from step 1, it might likely yield low accuracy. (Note that the labels of testing data are only used for evaluation.) In step 3, the trained model then runs inference on the unlabeled target domain training data, where the output predictions are kept as pseudo-labels for them. Next, in step 4, we take pairs of one source domain data and one target training data and then mix each pair of them as one semi-labeled training data. The mixing algorithm could be customized, such as WixUp tailored for wireless sensing data. Finally, in step 5, the new training data further fine-tunes the trained model. Thus, the model performance on the same testing data in the target domain should improve.
@@ 5 Experiments
  • First, we benchmark how WixUp enhances model performance in supervised learning across a variety of use cases, no matter what you use for datasets, model architectures, tasks, and sensing modalities. In general, WixUp helps increase the size of training data, which constantly provides a significant margin of improvement over no augmentation and outperforms other baseline augmentation methods.
  • The second group of experiments illustrates WixUp’s capability to significantly reduce the need for labeling data by utilizing unsupervised domain adaptation through self-learning.
  • Finally, we ablate study the algorithmic modules and hyper-parameters in our method to verify their effectiveness.
The following subsections begin with an elaboration on the experiment setup; then we present and analyze the results for these three groups of experiments.
5.1 Experiment setup
@@
5.1.1 Overall evaluation metrics
@@
5.1.2 Task-specific evaluation metrics
@@
To demonstrate the generalizability of our DA across different applications, we have selected three of the most common tasks in human tracking: keypoint pose estimation, action recognition, and user identification. For each task, we train and evaluate the model with WixUp to confirm that it reliably contributes to improvements in the evaluation data. This underscores WixUp’ usability in diverse applications.
@@
The evaluation metric for each task is 1) Mean Per Joint Position Error (MPJPE) for keypoint pose estimation and 2) classification accuracy for action recognition or user identification. In detail, MPJPE is one of the most common metrics for 3D human pose estimation or hand pose estimation. It calculates the average Euclidean distance between the predicted and the ground truth positions across all joints. MiliPoint uses MLE, so we also follow this usage in our paper. Besides, classification accuracy measures the proportion of total predictions that a model gets right, expressed in percentage. The goal of action recognition is to classify into one of the predefined actions (e.g., waving hands, limb extension, jumping up, etc.). For user identification, the objective is to recognize a user based on their behavioral or physical traits hidden in the detected sensor data, like movement patterns or interaction styles with devices.
5.1.3 Baseline augmentation methods
Three baselines serve as a comparison with WixUp to show its effectiveness under the same experimental settings: no augmentation, conventional global augmentation, and stacking. We run them in all the following benchmarking experiments.

### v2

@@ 3. Background: Wireless sensing and data formats
Most wireless sensing systems, including the standard mmWave radar (lovescu2020fundamentals, ), and many acoustic sensing and WiFi sensing systems use an FMCW radar modulation (nandakumar2015contactless, ; li2022lasense, ; zhao2018rf, ; zhao2018through, ) and hence our paper mainly focuses on this pipeline. Fig. 2 shows the processing pipeline of FMCW-based wireless sensing systems, which includes de-chirping and extraction of range, angle, velocity, and then 3D Cartesian coordinates. In detail, first, the transmitter (TX) emits chirps, which reflect off the subject and are then captured by the receiver (RX). These signals are processed through a mixer and filter to produce an Intermediate Frequency (IF) signal (or via algorithms outside the hardware if no mixer is available); A Fast Fourier Transform (FFT) is applied to measure range, followed by Doppler-FFT of the phases along the slow time axis (window axis) to measure velocity. The multi-channel phases can further estimate the angles of arrival. Then a Constant False Alarm Rate (CFAR) filtered the outputs by the noise level. Finally, the ranges and angles can be transformed to Cartesian coordinates along with velocity and signal intensity as 5D times series for downstream applications. Between modalities, the primary distinction lies in signal frequency. Different modalities’ data exhibits similar characteristics at range profile level. This pipeline enables detecting position and movement in human tracking or other applications.
@@
The challenge is the prevalent lack of raw data availability with the majority of public datasets, while the raw data, rich in detail, is critical for performing meaningful and effective DA. These datasets typically provide only Cartesian coordinates, i.e. the culmination of the data processing sequence. The issue with Cartesian coordinates is their high-level abstraction, owing to the filtering by CFAR, which leads to a considerable loss of fine-grained information that is imperative for in-depth DA. Moreover, the sparsity issue in the processed point clouds also makes it worse, while the range profile is a relatively dense representation. Unfortunately, it is intractable to reverse the CFAR algorithmically like inverse FFT.
@@
Hence, our goal is to propose a framework that can transform any data format of wireless signals into range profiles. This is because, unlike Cartesian coordinates, range profile is lossless, and is more interpretable than the raw data, allowing in-depth DA. The following subsections will detail how we achieve this goal, focusing on developing a way to inverse Cartesian coordinates to a simulated range profile.
4. System Design
Figure 3. A demonstration of the data mixing pipeline of WixUp.
@@ 4. System Design
As illustrated in Fig. 3, WixUp uses a common process of mixing-based DA. So, this section first provides a formal definition of the mix operation in the context of wireless signals, then details the customizations for mixing wireless signals at the range-profile level, and how it is embedded into the common mixing-based process. In brief, as in Fig 3, we first describe our solution to transform coordinates into range profiles using Gaussian mixtures, followed by a fault-tolerant mixing method leveraging the intersection of the Gaussians, with only O(n) computation complexity. Bootstrapping can further densify the mixed output. Finally, if the downstream model input is not a range profile but coordinates, we derive them along with the angles based on probability distribution. Furthermore, to test WixUp’s capability of reducing labeling efforts via self-learning, this section also introduces the self-training process that guides our experiment design.
In brief, this pipeline solves the challenges of the heterogeneous data availability of datasets, considering even the irreversible parts. And it is the first mixing DA algorithm customized for wireless signal characteristics.
4.1. Mixing-based data augmentation
@@
A typical mixing-based DA has inputs involving two sets of data: let $m$ and $n$ denote the sensing data and their ground truth. The input for a mixing operation is a pair of data $(m_{0},n_{0})$ and $(m_{1},n_{1})$. The output is synthesized data, $(m_{0}^{\prime},n_{0}^{\prime})$, in the same format as the two input tuples. Therefore, by iteratively mixing pairs of $(m_{i},n_{i})$ and $(m_{i+1},n_{i+1})$ where $i=0,1,...l$ ($l$ is the size of the original data), we could yield $l-1$ of synthesized data by one pass of iteration. Moreover, the distance $d$ between each pair can vary from 1 to 2 or more, i.e. $(m_{i},n_{i})$ and $(m_{i+d},n_{i+d})$. Thereby, by iterating with $d=1,2,...s$, we scale up the data by $s$ times.
@@
DA focuses on mixing $(m_{0},n_{0})$ and $(m_{1},n_{1})$, especially $m_{0}$ and $m_{1}$. Typically, $n$ is usually mixed by a convex combination, i.e. weighted average. In our work, we take the average, which is agnostic to task-specific ground truth format. For example, in pose estimation, $n$ could be a $19\times 3$ matrix, representing 19 body keypoints’ coordinates in 3D; thus averaging $n_{0}$ and $n_{1}$ equals the 3D skeleton located in the midway between the two original ones. Mixing is joint-wise so that it will mostly result in feasible positions, if the two original positions are not significantly distant within the same time sequence and the frame rate is sufficient. If these preconditions about frame rate and distance are not satisfied, infeasible positions such as shortened or intersected arms might occur. However, supervised learning use cases can mostly meet the conditions. In other cases, solutions such as filtering can be applied to avoid this issue. We will discuss this in detail in Sec. 6 based on our results and theory explanation. In tasks like user identification or action recognition, $n$ could be class ID, usually one-hot encoded. So, the average is a probability vector, which has an equal probability of distinct classes, or is identical if classes are the same. As for the data $m$, traditional image-based DA uses convex combination for pixel matrices similar to $n$. However, in wireless perception, $m$ could be raw signals, range profiles, 3D point clouds, or 5D time series of $(x,y,z,D,I)$ where $D$ is the Doppler velocity and $I$ is the signal intensity. As discussed above, directly adding them does not meaningfully align with the feature of wireless signals. Instead, WixUp aims to transform them all into the range profile level for mixing with lossless information. Next, we will introduce mixing $m_{0}$ and $m_{1}$ via a signal processing pipeline and probability-based algorithms.
4.2. WixUp data mixing pipeline
@@
4.2.1. Inverse Cartesian coordinates to simulated range profile using Gaussian mixture
@@
Most states in the FMCW data processing pipeline could freely transform to range profile using the standard signal processing algorithms, since they are mostly reversible, such as inverse FFT. However, the CFAR presents a unique challenge as it inherently discards a significant quantity of data, making reversal complex. Therefore, we put forth a novel inverse data processing pipeline that operates between the range data and Cartesian coordinates. This enables the application of WixUp to any public datasets, irrespective of the data format they are provided in, allowing us to bypass the limitation imposed by the absence of raw data.
@@
The core idea is to simulate the range profile as a Gaussian mixture model of Euclidean distances, as in the left of Fig. 3. In detail, first, we translate Cartesian coordinates to spherical coordinates as a representation better aligned with the transmission features of wireless signals. Next, dividing the Euclidean range by range resolution yields indices of bins, which are the means in a Gaussian mixture model; the probability density function (PDF) is derived as the range profile, i.e. the statistical distribution of the ranges. If the hyperparameters of the sensing system have a window size of 512 for de-chirping, and a range resolution of 3.75cm. Then the maximum allowable detection range is 19.2m. Assuming the human joint size is around 10-15cm, we use the standard Gaussian, implying its main lobe falls within the joint size.
Besides, choosing the low-level range profile as the mixing level means that it is also upward compatible with datasets with raw signal available or non-FMCW based algorithms. For instance, other features that can derive from parts of the standard pipeline could also work with our DA, such as Doppler velocity, Synthetic Aperture Radar (SAR) and Channel State Information (CSI). This work focuses and further optimizes on human tracking tasks using this standard pipeline, which is currently limited by the availability of open source large wireless datasets for benchmarking.
4.2.2. Mixing range profiles using intersection
@@
As illustrated in the center of Fig. 3, each pair of peaks may intersect midway along the range axis; the height of the intersected side lobe is roughly inversely proportional to the distance between the peaks. This approach involves taking middle points from all possible bisections of two sets of real points, as it is unclear which body joint each real point corresponds to; thus, we cannot restrict middle points to be extracted from identical joints. By considering all possible bisections, we ensure coverage of eligible synthetic points and introduce some ”temperature” into the process, borrowed from sampling in VAE. This makes the mixing fault-tolerant.
Instead of employing the closed-form algorithms for the intersection of Gaussian, we utilize an O(n) function to identify intersections. The validity of intersections can be determined using the formula $(b_{i}-a_{i})(b_{i-1}-a_{i-1})<0$, where $a$ and $b$ are the two PDF arrays and $(a_{i},a_{i-1})$ denotes a neighbor pair. This ensures that $ab$ pairs like $(0,0)(0,0)$ and $(0,1)(0,0)$ are excluded. This exclusion is crucial as such pairs do not represent valid intersection points. By implementing this method, we effectively filter out non-intersecting segments by one pass. This streamlines the intersection detection process, enhancing both efficiency and accuracy and does not rely on the complexity of the input Gaussian mixture.
The theory behind this mixing-based DA is the convex combination discussed in  (zhang2017mixup, ). Mixing is directly at the pixel level of images without manually designing features at any higher levels of representation, although it is less comprehensive. Similarly, in the context of wireless data, the physical representations of samples, as well as the environmental factors of sensors such as position and direction, are entangled representations in range profile. In other words, they have unified representations in range regardless of discrepancies in physical settings.
4.2.3. Boostrapping high-probability synthesis
Besides, we further bootstrap each intersection point weighting on its height. Since the height of the intersection is proportional to its validity as an eligible point, we use it as a weight to resample points around it. This follows the Monte Carlo simulation, where random samples are generated to simulate the properties of a system or to approximate integrals and sums. Besides, to add more randomness, we also keep the original peaks in the simulated range profile. They are assigned the lowest weight of one and all the weights of the above intersections are elevated by one as well. This step plus the above add-on algorithms in the mixing pipeline helps further improve the fidelity of augmented data. An exemplary fidelity is visualized in Fig. 3. Owing to the original sparsity nature of wireless signals as shown in Fig. 2, visualization can hardly justify the improvements in fidelity. So, we conduct comprehensive benchmark experiments to show its efficacy.
@@ 4.2.3. Boostrapping high-probability synthesis
4.2.4. Enrich range profile with angles based on probability distribution double-paned
@@
4.3. Use WixUp for unsupervised domain adaptation via self-training
We not only aim to evaluate WixUp in standard supervised learning, but also demonstrate its capability for reducing labeling efforts via self-training, achieved through our mixing scheme to generate new data. Note that non-mixing methods like random scaling are not applicable since they augment one data at a time, unlike mixing two data from different domains to close the domain gap. Mixing-based augmentation enables the creation of more realistic semi-real labels by mixing real data and predicted data, facilitating efficient domain adaptation through self-training.
Source domain & Target domain: The process is to adapt models trained on a source domain to inference well on a target domain. In this approach, the target domain dataset might have scarce labeled data or even no labeled data. In step 1 of this approach, the model is first trained with labeled data from the source domain. In step 2, partial target domain data is split out as training data and the rest is for testing. If we make predictions on testing target data using the trained model from step 1, it might likely yield low accuracy. (Note that the labels of testing data are only used for evaluation.) In step 3, the trained model then runs inference on the unlabeled target domain training data, where the output predictions are kept as pseudo-labels for them. Next, in step 4, we take pairs of one source domain data and one target training data and then mix each pair of them as one semi-labeled training data. The mixing algorithm could be customized, such as WixUp tailored for wireless sensing data. Finally, in step 5, the new training data further fine-tunes the trained model. Thus, the model performance on the same testing data in the target domain should improve.
@@ 5. Experiments
  • We benchmark how WixUp enhances model performance in supervised learning across a variety of use cases, no matter what you use for datasets, model architectures, tasks, and sensing modalities. In general, WixUp helps increase the size of training data, which constantly provides a significant margin of improvement over no augmentation and outperforms other baseline augmentation methods.
  • The second group of experiments illustrates WixUp’s capability to significantly reduce the need for labeling data by utilizing unsupervised domain adaptation through self-training.
  • We run ablation studies to verify the effectiveness of the algorithmic modules and hyper-parameters in WixUp.
The following subsections begin with an elaboration on the experiment setup, followed by results analysis for these three groups of experiments.
5.1. Experiment setup
@@
5.1.1. Overall evaluation metrics
@@
5.1.2. Task-specific evaluation metrics
@@
To demonstrate the generalizability of our DA across different applications, we selected three common tasks in human tracking: keypoint pose estimation, action recognition, and user identification. For each task, we train the model with WixUp to confirm that it reliably contributes to improvements in the evaluation data, highlighting WixUp’ usability in diverse applications.
@@
The evaluation metric for each task is 1) Mean Per Joint Position Error (MPJPE) for keypoint pose estimation and 2) classification accuracy for action recognition or user identification. In detail, MPJPE is common for 3D human pose estimation or hand pose estimation, calculating the average Euclidean distance between the predicted and the ground truth positions across key joints. MiliPoint uses mean localization error (MLE), so we also follow this usage in our paper. Besides, classification accuracy measures the percentage of correct predictions. The goal of action recognition is to classify into one of the predefined actions (e.g., waving hands, jumping up, etc.). User identification also uses accuracy, whose objective is to recognize a user based on their behavioral or physical traits hidden in the sensor data.
5.1.3. Baseline augmentation methods
We compare WixUp with three baselines, no augmentation, conventional global augmentation, and stacking, across all benchmark experiments to demonstrate its effectiveness. Other baselines that come with the original dataset work, such as shifting, are also listed in the following. Besides, temporal signal data augmentation, like stretching, is theoretically equivalent to shifting in the range profile that is included in our paper. Other traditional temporal signal augmentations are not applicable baselines since most of the open-source wireless datasets do not provide full raw data in the time domain. Finally, we also have a section comparing the existing DA methods in the wireless domain, although they are confined to specific tasks, models, or datasets, thus lacking a comprehensive benchmark evaluation.